The following Information applies to the questions displayed below. Delph Company uses a job-order costing system and has two manufacturing departments-Molding and Fabrication. The company provided the following estimates at the beginning of the year Mokaaion Total Machine-hours Fixed manufacturing overhead costs Variable manufacturing overhead per machine-hour 26,000 36,000 62,000 250.000 $1.020,000 770.000 s 5.70 5.70During the year, the company had no beginning or ending Inventories andt started, completed, and sold only two jobs-J00 D-70 and Job C-200. tt provided the following information related to those two jobs: Job D-70: Molding Fabrication Total Direct materias s3s 328,000699,000 cost Direct labor 220,000 $140,000 s360,000 cost Machine- 18,000 8,000 26,000 hours Job C-200: Moiding Fabrication Total Direct 270,000250,000 520,000 cost Direct labor s 150,000 s 250,000 s 400,000 cost Machine- 8,000 28,000 36,000 hours Delph had no overapplied or underapplied manufacturing overhead during the yeac value: 3.00 points Assume Delph uses a plantwide overhead rate based on machine-hours. Required: 1- Compute the predetermined plantwide overhead rate. (Round your a. answer to 2 decimal places.) per MH 1- Compute the total manufacturing costs assigned to Job D-70 and Job C- b. 200. (Round your intermediate calculations to 2 decimal places.) Job D-70 Job C- 200 1- If Delph establishes bid prices that are 140% of total manufacturing costs. c. what bid price would it have established for Job D-70 and Job C-200? Round your intermediate calculations to 2 decimal places.) Job D-70 Job C-200 1- What is Delphs cost of goods sold for the year? (Round your d. intermediate calculations to 2 decimal places.)
Answer #1

Part 1 – A

Predetermined overhead rate = variable costs per machine hour + fixed costs per machine hour = 5.70+(1020000/62000) = $22.15 per MH

Fixed cost per machine hour = fixed manufacturing overhead costs/machine hours

Therefore,

Predetermined overhead rate

$22.15

Per MH

Part 1 B

Job D-70

Job C-200

Direct materials cost

699000

520000

Direct labor cost

360000

400000

Manufacturing overhead

575900 (26000*22.15)

797400 (36000*22.15)

Total manufacturing costs

1634900

1717400

Therefore,

Total manufacturing costs

Job D-70

1634900

Job C-200

1717400

Part 1 – C

Job D-70

Job C-200

Total manufacturing costs

1634900

1717400

margin

140%

140%

Bid price

2288860

2404360

Therefore,

Bid price

Job D-70

2288860

Job C-200

2404360

Part 1 D

Cost of goods sold = total manufacturing cost for Job D-70 + total manufacturing cost for Job C-200 = 1634900+1717400 = 3352300

Therefore,

Cost of goods sold

$3352300